a. What are the confidence limits (α=0.05) for the difference between the true mean value of Y when X1 = 3 and the true mean value of Y when X2 = -2?

Fe (X) Loss in MDD (Y)

 1  0.01 127.6
 2  0.48 124.0
 3  0.71 110.8
 4  0.95 103.9
 5  1.19 101.5
 6  0.01 130.1
 7  0.48 122.0
 8  1.44  92.3
 9  0.71 113.1
10  1.96  83.7
11  0.01 128.0
12  1.44  91.4
13  1.96  86.2

Thirteen specimens of 90/10 Cu-Ni alloys, each with a specific iron were tested in a corrosion-wheel setup. The wheel was rotated in salt seawater at 30 ft/s for 60 days. The corrosion was measured in weight loss in milligram/square decimeter/day, MDD. The following data were collected. Show all work.
Answer #1

a) The 95% Confidence interim for the contrast between evident mean estimations of Y for X1 = 3 and X2 = - 2 can be effectively determined by the recipe:

where Y1 & Y2 are the anticipated mean qualities of Y for X1 = 3 & X2 = - 2 respectively. here i have done some calculations you can find the values by substituting in above formula

Sample Size , n=   13, t critical value=   t α/2 =    2.20098516 ,SSE=   (Sx*Sy - S²xy)/Sx =    102.85,

Se =    √(SSE/(n-2)) =    3.0578,

slope, B1 SSxy/SSxx-24.01989345 intercept, β0-y-B1*x 129.7865993

X Value=   3
Confidence Level=   95%
  
Sample Size , n=   13
Degrees of Freedom,df=n-2 =   11
critical t Value=tα/2 =   2.201
X̅ =Σx/n = 0.873
Σ(x-x̅)² =Sxx   5.709
Standard Error of the Estimate,Se=   3.058
h Statistic = (1/n+(X-X̅)^2/Sxx) =    0.869
Predicted Y (YHat) =
Ŷ=   129.7866 -24.0199   *3 = 57.727

(y y xy y) 0.01 127.60.7449018 352.8618 -16.2126 0.48 124 0.1545095 230.5725-5.96872 3.938698 -0.32364 0.95 103.9 0.0059172 24.16101 -0.37811 .51485 -2.31841 0.01 130.10.7449018 453.0349 -18.3703 0.48 122 0.1545095 173.8341-5.18257 0.3214018 272.75799.36295 0.71 113.10.0265941 18.35793 -0.69872 630.7825 -27.2985 0.01 128 0.7449018 368.0495 -16.5578 1.44 91.4 0.3214018 303.2956 -9.87318 1.96 86.2 1.1814018 511.4556 -24.5812 (x x^ 0.71 110.8 0.0265941 1.19 101.5 0.1004402 53 1.44 92.3 196 83.711814018

total sum 11.35 1414.6 5.708877 3396.617-137.127 0.873076923 108.815385SSxxSSyy SSxy